answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability that a family of 4 would have a girl first and a girl last?

There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.

A family of 4 is a family of two parents and two children. The probability that both children are girls is 0.2334

What is the probability of a set of 28 people having two sets of matched birthdays?

Prob(two sets of matched birthdays)

= 1 - [Prob(No matched birthdays)] + [Prob(One set of matched birthdays)]

= 0.63

What is the probability of A and B telling lie Prob of a telling truth equals 80 percent Prob of b telling truth equals 75 percent?

Chance that A told the truth: 0.8 (80%)

Chance that B told the truth: 0.75 (75%)

Chance that A told a lie: 1 - 0.8 = 0.2 (20%)

Chance that B told a lie: 1 - 0.75 = 0.25 (25%)

Chance that both A and B told a lie: 0.2 * 0.25 = 0.05 (5%)

Chance that both A and B told the truth: 0.8 * 0.75 = 0.6 (60%)

Chance that A told a lie and B told the truth: 0.2 * 0.75 = 0.15 (15%)

Chance that A told the truth and B told a lie: 0.8 * 0.25 = 0.2 (20%)

How many standard deviations are 95 percent of measurements away from the mean?

95 percent of measurements are less than 2 standard deviations away from the mean, assuming a normal distribution.

What is the probability that a recessive trait will appear?

It depends on the genotype of the parents, but assuming there is an equal chance of being dominant homozygous, recessive homozygous or heterozgous and there are only two possible genes, there is a 1 in 4 chance that the recessive trait will appear.

How many possible outcomes for one spin and one coin toss?

Two times the number of outcomes of the spin - which is not specified in the question.

What is the probability of picking are red card from a deck of cards?

Since half of the cards (26 of the 52) are a red card (hearts and diamonds) the odds are 1 in 2 (or 50%) that any given card will be red.

What event in?

Recreational Basketball and soon Baseball. Young Children 6 and 8 years old. Mother previously showed no interest when family was together never even attended practices or games. Now trying to keep family and friends she doesn't like from attending. No one speaks or harasses her and she claims she will be filing charges against my friends and family for attending and making her FEEL harassed.

What is the probability that a random person has a birthday in July?

A simple answer is 1/12 = 0.0833... since there are 12 months in a year, of which 1 has been selected.

A slightly better estimate, allowing for the the fact that July has 31 days is 31/365 = 0.0849. Adjusting for leap years does not affect the probability at 4 decimal places.

These answers assume that births are uniformly distributed over the year. In fact they are not but comprehensive information on births by month is not readily available.

What is the probability of rolling an odd number on die?

The probability of rolling a odd number on a standard die is 3 in 6, or 1 in 2, or 0.5.
You don't say with what, so I assume with a standard six sided die. The probability of rolling an odd number with a standard six sided die is 3 in 6, or 1 in 2, or 0.5.

Find the probability that in 5 tosses of a fair die a 3 appears at no one time?

Possible outcomes per each toss = 6

Number of them that are not a 3 = 5

Probability of no 3 in one toss = 5/6.

Probability of no 3 in 5 tosses = (5/6) x (5/6) x (5/6) x (5/6) x (5/6) = 3,125/7,776 = 0.40188 = 40.19%(rounded)

10 persons can watch a movie in 7 days What is the probability that at least two them watch on the same day?

Not sure but could it be:-

Probability of selecting one day out of seven:-(1/7)

Now probability that only one person out of 10 watches the movie:-(1/10)

Thus the total probability of only one person watching the movie on any given day will be:- (1/7)*(1/10)=1/70

Now, the probability that at least two of them watch the movie on the same day will be:- 1-(1/70)=(69/70) Hope this works, Any other suggestions are welcome.

- Regards, NightHawk